OpenStudy (anonymous):

What is the speed, in meters per second, of a paper airplane that flies 24 meters in 6 seconds? A. 144 B. 30 C. 18 D. 4 What is the speed, in meters per second, of a paper airplane that flies 24 meters in 6 seconds? A. 144 B. 30 C. 18 D. 4 @Mathematics

OpenStudy (amistre64):

spose you go 60 miles in 2 hours; then that means that youve gone 30 miles every hour. 60/2 = 30 distance/time = speed

OpenStudy (anonymous):

thanks bro

OpenStudy (amistre64):

i just noticed that I should have used different numbers in my example. the process is good, but i shouldnt have used a result that can be confused with the actual answer. 24 is the speed in your problem, and 6 is the time. s = d/t s = 24/6 = 4 is the specific answer to this question.